OpenStudy (dejuan1):

You place $4,000.00 in a bank account with an interest rate of 5.25% APR and another $2,000.00 in an account with an interest rate of 6.00% APR. After 6 months, what is the difference in interest that the accounts earn

jimthompson5910 (jim_thompson5910):

The formula you'll use is the compound interest formula \[\LARGE A = P*\left(1+\frac{r}{n}\right)^{n*t}\] A = final amount in account after t years P = amount deposited r = interest rate n = compound frequency (how many times you compound interest per year) t = time in years

jimthompson5910 (jim_thompson5910):

keep in mind that t in years so 6 months = 6/12 = 1/2 = 0.5 years

OpenStudy (dejuan1):

thanks

jimthompson5910 (jim_thompson5910):

no problem